OpenStudy (anonymous):

HELP PLEASE! MULTIPLE CHOICE! Write the equations of the asymptotes for the hyperbola represented by 36x^2-y^2-4y=-32.

OpenStudy (anonymous):

a. y - 1 = +/- 6 (x - 2) b. y = +/- 6x c. y + 2 = +/- 6(x - 1) d. y + 2 = +/- 6x

OpenStudy (anonymous):

@Destinymasha @Loser66

OpenStudy (anonymous):

Help me please! I dont understand...

OpenStudy (loser66):

first off, you have to write the expression under the standard form of hyperbola.

OpenStudy (loser66):

which is \[\dfrac{x^2}{a^2}-\dfrac{y^2}{b^2}=1\] then, apply the formula to find out the asymptote which is \[y =\pm\dfrac{-b}{a}\]

OpenStudy (anonymous):

Standard form: (((y+2)^2)/(6^2)) - ((x^2)/(1^2)) = 1 @Loser66

OpenStudy (anonymous):

Is that right?

OpenStudy (loser66):

Let me check, honestly I didn't work it out. just gave you steps. :)

OpenStudy (anonymous):

Ok

OpenStudy (loser66):

yup

OpenStudy (loser66):

but your hyperbola is another form, so that the asymptote will be \(y =\pm \dfrac{-a}{b}\)where a is the number under "y" term and b is the number under "x" term. Got what I mean?

OpenStudy (loser66):

I am sorry, \(y = \pm\dfrac{-a}{b}x\)

OpenStudy (anonymous):

Ok so the answer is b? @Loser66

OpenStudy (loser66):

yup